derive the slope-intercept form of a linear equation from the slope formula using (o,b) as the coordinates of the point where the line crosses the y-axis and arbitrary point (x,y).
Answer by stanbon(75887)

derive the slope-intercept form of a linear equation from the slope formula using (o,b) as the coordinates of the point where the line crosses the y-axis and arbitrary point (x,y).
---------------
You are given two points: (0,b) and (x,y)
slope = change in y / change in x = y-b/x-0 = (y-b)/x
So, m=(y-b)/x
Solve for y as follows:
mx = y-b
y = mx+b
==========
Cheers,
Stan H.